OpenStudy (anonymous):

I need HELP!!! please! Solve. Write each answer in simplest form. 1. a-1/2 = 1/4 2. m + 1/6 = 5/6 3. p - 2/3 =5/6 4. 1/5x = 8 5. 2/3r = 3/5 6. 3w = 3/7 8. n - 1/5 = 3/5 9. t - 3/8 = 1/4 10. s - 7/24 = 1/3 11. x + 2/3 = 2/7/8 12. h + 7/10 = 7/10 13. y + 5/6 = 19/20 14. 1/5x = 4 15. 1/4w = 1/8 16. 5y = 3/20 17. 6z = 1/2 18. 5/8x = 2/5 19. 5/8n = 1/1/5 (I skipped 7 for a reason lol) please help!

OpenStudy (anonymous):

For a reason what you mean lol

OpenStudy (anonymous):

a-1/2 = 1/4=3/4

OpenStudy (anonymous):

m + 1/6 = 5/6=m=2/3

OpenStudy (anonymous):

p - 2/3 =5/6=3/2

OpenStudy (anonymous):

1/5x = 8=1/40

OpenStudy (anonymous):

. 2/3r = 3/5=10/9

hero (hero):

Wow, that's a lot of problems

OpenStudy (anonymous):

True a lot

OpenStudy (anonymous):

ikr?? and thanks sooooooo much for helping guys i dont understand any of this my teacher just lays work on us n doesnt even explain it

OpenStudy (anonymous):

3w = 3/7=1/7

OpenStudy (anonymous):

Guys? I think I am the only one helping so why are you thanking like guys lolz

hero (hero):

That's a recurring theme in grammar school high school. If you ever become a math major, the recurring theme will be "I understand math, but not proofs". lol

hero (hero):

I would help but I'm busy

OpenStudy (ash2326):

@Makayla340 understand one problem and you'll be able to solve all:)

OpenStudy (anonymous):

n - 1/5 = 3/5 =4/5

hero (hero):

^ What ash said

OpenStudy (anonymous):

give a man a fish teach a man to fish ...

OpenStudy (anonymous):

@ash2326 i dont know any of it like i said my teacher gives us like 5 sections to do n never explains it! -.-

OpenStudy (anonymous):

but i understand what you guys r saying

OpenStudy (anonymous):

Then tell your principal

hero (hero):

All you have to do is just isolate the variable. If you already know how to add and subtract fractions, this is easy.

OpenStudy (anonymous):

@kumar2006 she's even worse...

OpenStudy (anonymous):

Change your school

OpenStudy (ash2326):

Ok pick any of these, I'll illustrate you and then you'd try the rest.:)

hero (hero):

How do you contact the principal if you do online school? lol

OpenStudy (anonymous):

I dont do online school lol

OpenStudy (anonymous):

and okay @ash2326

OpenStudy (anonymous):

Give me medals I helped you on a lot =)

OpenStudy (anonymous):

Done :) and thanksss!!

OpenStudy (anonymous):

No doubt

OpenStudy (ash2326):

Then tell , which one of these you want me to explain

OpenStudy (anonymous):

agreed

OpenStudy (ash2326):

Which one Makayla???

OpenStudy (anonymous):

Where you from Makayla?

OpenStudy (anonymous):

5/8x = 2/5 @ash2326 and @kumar2006 St. thomas United States Virgin Islands

OpenStudy (anonymous):

Oh great, Nice place

OpenStudy (anonymous):

You've heard of it?

OpenStudy (anonymous):

Most people haven't.

OpenStudy (anonymous):

i havenet

OpenStudy (anonymous):

Lol, that's funny... but yeah it is! My family originates from many places so I'm mixed

OpenStudy (ash2326):

We have \[\frac{5}{8} x=\frac{2}{5}\] We have to get on one side just x, with coefficient 1 and then the right hand side's value , will be the solution. Let's see how we can get just x multiply both sides by 8 \[8 \times \frac{5}{8} x=8 \times\frac{2}{5}\] Now we have \[5x=\frac{16}{5}\] Now divide both sides by 5 \[\frac{5x}{5}=\frac{16}{5\times 5}\] we get \[x=\frac{16}{25}\]

OpenStudy (anonymous):

What's coefficient 1? He didnt tell us what it is! ughh im stressed noww

OpenStudy (ash2326):

suppose I have 2x. here coefficient of x is 2 if I have 1x or x then the coefficient is 1

OpenStudy (anonymous):

ooh okaii is there like a certain way you find a coefficient or is it like an automatic find

OpenStudy (anonymous):

Are you also west indian?

OpenStudy (anonymous):

me?

OpenStudy (anonymous):

Yes yo

OpenStudy (anonymous):

uhh yeah but im also irish dutch russian and native american lol

OpenStudy (anonymous):

Wow ever been to any pow wow?

OpenStudy (anonymous):

wat is that?

OpenStudy (anonymous):

Like native american festival lolz

OpenStudy (anonymous):

oh noo i havent gotten a chance to but i would like to

OpenStudy (anonymous):

Lolz

OpenStudy (ash2326):

It's just the numeral along with x

OpenStudy (anonymous):

ahhh okay i sorta have an idea

OpenStudy (anonymous):

thanks so much ash :)

OpenStudy (anonymous):

Ash is a life saver

OpenStudy (anonymous):

IKR?? and you too thanks! :)

OpenStudy (anonymous):

I didn't do anything

OpenStudy (anonymous):

yeah you did lol you answered the questions for me lol

OpenStudy (anonymous):

Lolz really bahahahah just kidding

OpenStudy (anonymous):

lmbo XD nice one

OpenStudy (anonymous):

In what country do you guys live?

OpenStudy (anonymous):

I am from Nyc

OpenStudy (ash2326):

Welcome @Makayla340 and thanks @kumar2006

OpenStudy (anonymous):

so you live there?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

mhm awesomesauce ;) lol !

OpenStudy (anonymous):

Sauce?

OpenStudy (anonymous):

Ever been here

OpenStudy (anonymous):

lol it's a word my brother n i made up >.<

OpenStudy (anonymous):

Yes! I have family there

OpenStudy (anonymous):

but i dont remember which part

OpenStudy (anonymous):

Really where exactly so I can go meet them lmfaoo j.k

OpenStudy (anonymous):

lollll!!!!!!! i went there when i was like six so it was a while agoo but i remember n have oictures good thing we went in summer time

OpenStudy (anonymous):

pictures*

OpenStudy (anonymous):

Really wow that was long ago.

OpenStudy (anonymous):

yerp :p